# Name Tagging Under Domain Shift via Metric Learning for Life Sciences

### v1

@@ Appendix G Scientific Artifacts
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is directed to or away from a particular cellular component <AtLoc><FromLoc><ToLoc> or subcellular location within the cell.
Transport
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is moved or conveyed to or away from a particular cellular component <FromLoc><ToLoc> or subcellular location within the cell.
Gene Expression
@@
Indicated by the given trigger <Trigger>, genetic information from a gene <Theme> is used to produce a functional gene product, such as RNA or protein.
Transcription
@@
Indicated by the given trigger <Trigger>, genetic information from a gene <Theme> is transcribed into RNA, usually messenger RNA (mRNA), by RNA polymerase.
Translation
@@
Indicated by the given trigger <Trigger>, the genetic information carried by mRNA <Theme> is used to synthesize a protein by ribosomes in the cell.
Degradation
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes breakdown or degradation into smaller components.
Binding
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> interacts and forms a complex with another molecule(s) resulting in the product of a molecular complex <Product>.
@@
Type
Template
Binding
Indicated by the given trigger <Trigger>, a specific molecule <Theme> interacts and forms a complex with another molecule(s) resulting in the product of a molecular complex <Product>.
Dissociation
Indicated by the given trigger <Trigger>, a specific complex <Theme> breaks apart, resulting in the release of individual molecules <Product> as products.
Regulation
Indicated by the given trigger <Trigger>, an entity <Theme> is controlled or influenced by another entity <Cause> to achieve a specific biological effect or outcome.
Positive Regulation
Indicated by the given trigger <Trigger>, an entity <Theme> is promoted or enhanced by another entity <Cause> to achieve a specific biological effect or outcome.
Activation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is stimulated or facilitated by another entity <Cause> to increase its activity, function, or biological effect.
Negative Regulation
Indicated by the given trigger <Trigger>, an entity <Theme> is inhibited or suppressed by another entity <Cause> to achieve a specific biological effect or outcome.
Inactivation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is deactivated or rendered inactive by another entity <Cause>, leading to a reduction or cessation of its biological function.
Pathway
Indicated by the given trigger <Trigger>, involving one or more molecules <Participant> that collaborate to accomplish a specific biological function or response.
@@
Type
Template
Gene Expression
Indicated by the given trigger <Trigger>, a specific protein or a group of genes <Theme> are involved in the transcription and translation of genetic information to produce functional gene products, such as RNA or protein.
Transcription
Indicated by the given trigger <Trigger>, a specific protein or a group of genes <Theme> are involved in the synthesis of RNA from DNA template by RNA polymerase.
Protein Catabolism
Indicated by the given trigger <Trigger>, a specific protein <Theme> is broken down or degraded into smaller peptide fragments or amino acids.
Phosphorylation
Indicated by the given trigger <Trigger>, a specific protein <Theme> undergoes the addition of a phosphate group at a particular site <Site>, resulting in the modification of the protein’s structure and function.
Localization
Indicated by the given trigger <Trigger>, a specific core entity <Theme> is directed to or away one location <AtLoc><ToLoc> within the cell or organism.
Binding
Indicated by the given trigger <Trigger>, a specific core entity <Theme> interacts and forms a connection with another entity <Site>, leading to the formation of a complex or association.
Regulation
Indicated by the given trigger <Trigger>, a specific core entity or event <Theme> is controlled or influenced by another core entity or event <Cause> through interactions at specific sites on molecules or entities <Site><CSite>, potentially resulting in modulation of biological processes.
Positive Regulation
Indicated by the given trigger <Trigger>, a specific core entity or event <Theme> is promoted or enhanced by another core entity or event <Cause> through interactions at specific sites on molecules or entities <Site><CSite>, potentially resulting in an increase in the intensity or rate of a biological process.
Negative Regulation
Indicated by the given trigger <Trigger>, a specific core entity or event <Theme> is inhibited or suppressed by another core entity or event <Cause> through interactions at specific sites on molecules or entities <Site><CSite>, potentially resulting in a decrease in the intensity or rate of a biological process.
Process
Indicated by the given trigger <Trigger>, involving a core entity that collaborates to accomplish a specific biological function or response.
@@
Indicated by the given trigger <Trigger>, a specific anatomical or pathological entity <Theme> undergoes progressive changes or growth, leading to the formation of a more complex and specialized structure or condition over time.
Blood Vessel Development
@@
Indicated by the given trigger <Trigger>, blood vessels <Theme> undergo progressive changes or growth at a specific location <AtLoc>, leading to the formation and maturation of the vascular network.
Growth
@@
Indicated by the given trigger <Trigger>, a specific anatomical or pathological entity <Theme> undergoes an increase in size, quantity, or complexity over time.
Death
@@
Indicated by the given trigger <Trigger>, a specific anatomical or pathological entity <Theme> ceases to exhibit signs of life and undergoes irreversible loss of vital functions.
Cell Death
@@
Indicated by the given trigger <Trigger>, a specific cell <Theme> undergoes a series of events leading to its own demise, often through programmed cell death or other mechanisms.
Breakdown
@@
Indicated by the given trigger <Trigger>, a specific anatomical or pathological structure <Theme> disintegrates, decomposes, or undergoes degradation over time.
Cell Proliferation
@@
Indicated by the given trigger <Trigger>, a specific cell <Theme> undergoes rapid and controlled replication or division, leading to an increase in the number of daughter cells.
Cell Division
@@
Indicated by the given trigger <Trigger>, a specific cell <Theme> divides into two or more daughter cells through mitosis or meiosis.
Cell Differentiation
@@
Indicated by the given trigger <Trigger>, a specific cell <Theme> undergoes changes in gene expression and morphology to become specialized and acquire distinct functions at a specific anatomical or pathological location <AtLoc>.
Remodeling
@@
Indicated by the given trigger <Trigger>, a specific tissue <Theme> undergoes structural changes, reorganization, and modification in response to various stimuli or during growth and development.
Reproduction
@@
Indicated by the given trigger <Trigger>, a specific organism <Theme> produces offspring through sexual or asexual means, leading to the continuation of the species.
Mutation
@@
indicated by the given trigger <Trigger>, a specific gene, genome, or protein <Theme> undergoes a heritable change in its genetic sequence or structure at a particular site <Site>, potentially leading to alterations in its function or expression within a specific anatomical or pathological context <AtLoc>.
Carcinogenesis
@@
Indicated by the given trigger <Trigger>, a specific anatomical or pathological entity <Theme> undergoes a series of genetic and cellular changes at a specific anatomical or pathological location <AtLoc>, leading to the development of cancer.
Cell Transformation
@@
Indicated by the given trigger <Trigger>, a specific cell <Theme> undergoes changes in its phenotype, function, or behavior at a specific anatomical or pathological location <AtLoc>, often associated with the acquisition of abnormal or cancerous characteristics.
Metastasis
@@
Indicated by the given trigger <Trigger>, a specific anatomical or pathological entity <Theme> spreads and establishes secondary growths or lesions at a different anatomical or pathological location <ToLoc> from the primary tumor site.
Infection
@@
Indicated by the given trigger <Trigger>, an organism <Participant> invades and establishes itself in a specific anatomical or pathological site <Theme>, leading to a disease condition.
Metabolism
@@
Indicated by the given trigger <Trigger>, collective chemical reactions occur within an organism <Theme> involving the processing, transformation, and utilization of specific molecules to maintain cellular functions and energy requirements.
Synthesis
@@
Indicated by the given trigger <Trigger>, a specific simple chemical <Theme> is produced or created through chemical reactions or biological processes.
Catabolism
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes chemical reactions or metabolic pathways to break down into simpler compounds, releasing energy in the process.
Amino Acid Catabolism
@@
Indicated by the given trigger <Trigger>, a specific amino acid <Theme> is broken down through metabolic pathways, leading to the release of energy and the generation of byproducts like ammonia and carbon compounds.
Glycolysis
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes a series of enzymatic reactions, ultimately converting glucose into pyruvate and producing ATP and NADH as energy carriers.
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of carbohydrate molecules (glycans) to specific sites, typically on proteins or lipids, leading to the formation of glycoproteins or glycolipids with diverse biological functions.
Acetylation
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of an acetyl group at a particular site <Site>, catalyzed by another molecule <Cause>.
Deacetylation
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> has an acetyl group removed from a particular site <Site>, facilitated by another molecule <Cause>.
Ubiquitination
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is modified by the attachment of one or more ubiquitin molecules at a particular site, facilitated by another molecule <Cause>, often involving a simple chemical group <Site> as the site of attachment.
Deubiquitination
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> has ubiquitin molecules removed from a particular site, facilitated by another molecule <Cause> involving a simple chemical group <Site> as the site of removal.
Gene Expression
@@
Indicated by the given trigger <Trigger>, a specific gene, genome, or protein <Theme> is activated, leading to the production of RNA or protein and subsequent biological functions.
Transcription
@@
Indicated by the given trigger <Trigger>, genetic information from a specific gene, genome, or RNA molecule <Theme> is used as a template to produce complementary RNA (usually messenger RNA) by RNA polymerase.
Translation
@@
Indicated by the given trigger <Trigger>, the genetic information carried by messenger RNA <Theme> is used to synthesize a protein by ribosomes in the cell.
Protein Processing
@@
Indicated by the given trigger <Trigger>, the series of post-translational modifications, folding, and transportation of a specific gene, genome, or protein <Theme> in the cell to achieve its mature and functional form.
Phosphorylation
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of a phosphate group at a particular site <Site> within a protein domain or region, often regulating the molecule’s activity or function.
Dephosphorylation
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> has a phosphate group removed from a particular site <Site> within a protein domain or region, often resulting in the modulation or termination of its activity or function.
DNA Methylation
@@
Indicated by the given trigger <Trigger>, a specific gene or genome <Theme> has a methyl group added to a particular site <Site> within a protein domain or region, often resulting in the regulation of gene expression and epigenetic modifications.
DNA Demethylation
@@
Indicated by the given trigger <Trigger>, a specific gene or genome <Theme> has a methyl group removed from a particular site <Site> within a protein domain or region, often resulting in the modulation of gene expression and epigenetic modifications.
Pathway
@@
Indicated by the given trigger <Trigger>, involving a specific molecule <Participant> that collaborates to accomplish a specific biological function or response.
Binding
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> forms a physical association or interaction with another molecule or region <Site> on a protein or DNA, potentially leading to functional changes or regulatory effects.
Dissociation
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> separates or detaches from another molecule or region <Site> on a protein or DNA, leading to the termination or disruption of their interaction or complex formation.
Localization
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is directed to or away from a particular cellular component <AtLoc><FromLoc><ToLoc> or subcellular location within the cell.
Regulation
@@
Indicated by the given trigger <Trigger>, an entity <Theme> is controlled or influenced by another entity <Cause> to achieve a specific biological effect or outcome.
Positive Regulation
@@
Indicated by the given trigger <Trigger>, an entity <Theme> is promoted or enhanced by another entity <Cause> to achieve a specific biological effect or outcome.
Negative Regulation
@@
Indicated by the given trigger <Trigger>, an entity <Theme> is inhibited or suppressed by another entity <Cause> to achieve a specific biological effect or outcome.
Planned Process
@@
Indicated by the given trigger <Trigger>, a specific entity <Theme> is involved or manipulated using an instrument <Instrument> for a predetermined outcome or purpose.

### v2

@@ Appendix G Scientific Artifacts
Localization Indicated by the given trigger <Trigger>, a specific molecule <Theme> is directed to or away from a particular cellular component <AtLoc><FromLoc><ToLoc> or subcellular location within the cell.
@@
Transport Indicated by the given trigger <Trigger>, a specific molecule <Theme> is moved or conveyed to or away from a particular cellular component <FromLoc><ToLoc> or subcellular location within the cell.
@@
Gene Expression Indicated by the given trigger <Trigger>, genetic information from a gene <Theme> is used to produce a functional gene product, such as RNA or protein.
@@
Transcription Indicated by the given trigger <Trigger>, genetic information from a gene <Theme> is transcribed into RNA, usually messenger RNA (mRNA), by RNA polymerase.
@@
Translation Indicated by the given trigger <Trigger>, the genetic information carried by mRNA <Theme> is used to synthesize a protein by ribosomes in the cell.
@@
Degradation Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes breakdown or degradation into smaller components.
@@
Binding Indicated by the given trigger <Trigger>, a specific molecule <Theme> interacts and forms a complex with another molecule(s) resulting in the product of a molecular complex <Product>.
@@
Type Template
Binding Indicated by the given trigger <Trigger>, a specific molecule <Theme> interacts and forms a complex with another molecule(s) resulting in the product of a molecular complex <Product>.
Dissociation Indicated by the given trigger <Trigger>, a specific complex <Theme> breaks apart, resulting in the release of individual molecules <Product> as products.
Regulation Indicated by the given trigger <Trigger>, an entity <Theme> is controlled or influenced by another entity <Cause> to achieve a specific biological effect or outcome.
Positive Regulation Indicated by the given trigger <Trigger>, an entity <Theme> is promoted or enhanced by another entity <Cause> to achieve a specific biological effect or outcome.
Activation Indicated by the given trigger <Trigger>, a specific molecule <Theme> is stimulated or facilitated by another entity <Cause> to increase its activity, function, or biological effect.
Negative Regulation Indicated by the given trigger <Trigger>, an entity <Theme> is inhibited or suppressed by another entity <Cause> to achieve a specific biological effect or outcome.
Inactivation Indicated by the given trigger <Trigger>, a specific molecule <Theme> is deactivated or rendered inactive by another entity <Cause>, leading to a reduction or cessation of its biological function.
Pathway Indicated by the given trigger <Trigger>, involving one or more molecules <Participant> that collaborate to accomplish a specific biological function or response.
@@
Type Template
Gene Expression Indicated by the given trigger <Trigger>, a specific protein or a group of genes <Theme> are involved in the transcription and translation of genetic information to produce functional gene products, such as RNA or protein.
Transcription Indicated by the given trigger <Trigger>, a specific protein or a group of genes <Theme> are involved in the synthesis of RNA from DNA template by RNA polymerase.
Protein Catabolism Indicated by the given trigger <Trigger>, a specific protein <Theme> is broken down or degraded into smaller peptide fragments or amino acids.
Phosphorylation Indicated by the given trigger <Trigger>, a specific protein <Theme> undergoes the addition of a phosphate group at a particular site <Site>, resulting in the modification of the protein’s structure and function.
Localization Indicated by the given trigger <Trigger>, a specific core entity <Theme> is directed to or away one location <AtLoc><ToLoc> within the cell or organism.
Binding Indicated by the given trigger <Trigger>, a specific core entity <Theme> interacts and forms a connection with another entity <Site>, leading to the formation of a complex or association.
Regulation Indicated by the given trigger <Trigger>, a specific core entity or event <Theme> is controlled or influenced by another core entity or event <Cause> through interactions at specific sites on molecules or entities <Site><CSite>, potentially resulting in modulation of biological processes.
Positive Regulation Indicated by the given trigger <Trigger>, a specific core entity or event <Theme> is promoted or enhanced by another core entity or event <Cause> through interactions at specific sites on molecules or entities <Site><CSite>, potentially resulting in an increase in the intensity or rate of a biological process.
Negative Regulation Indicated by the given trigger <Trigger>, a specific core entity or event <Theme> is inhibited or suppressed by another core entity or event <Cause> through interactions at specific sites on molecules or entities <Site><CSite>, potentially resulting in a decrease in the intensity or rate of a biological process.
Process Indicated by the given trigger <Trigger>, involving a core entity that collaborates to accomplish a specific biological function or response.
@@
Development Indicated by the given trigger <Trigger>, a specific anatomical or pathological entity <Theme> undergoes progressive changes or growth, leading to the formation of a more complex and specialized structure or condition over time.
@@
Blood Vessel Development Indicated by the given trigger <Trigger>, blood vessels <Theme> undergo progressive changes or growth at a specific location <AtLoc>, leading to the formation and maturation of the vascular network.
@@
Growth Indicated by the given trigger <Trigger>, a specific anatomical or pathological entity <Theme> undergoes an increase in size, quantity, or complexity over time.
@@
Death Indicated by the given trigger <Trigger>, a specific anatomical or pathological entity <Theme> ceases to exhibit signs of life and undergoes irreversible loss of vital functions.
@@
Cell Death Indicated by the given trigger <Trigger>, a specific cell <Theme> undergoes a series of events leading to its own demise, often through programmed cell death or other mechanisms.
@@
Breakdown Indicated by the given trigger <Trigger>, a specific anatomical or pathological structure <Theme> disintegrates, decomposes, or undergoes degradation over time.
@@
Cell Proliferation Indicated by the given trigger <Trigger>, a specific cell <Theme> undergoes rapid and controlled replication or division, leading to an increase in the number of daughter cells.
@@
Cell Division Indicated by the given trigger <Trigger>, a specific cell <Theme> divides into two or more daughter cells through mitosis or meiosis.
@@
Cell Differentiation Indicated by the given trigger <Trigger>, a specific cell <Theme> undergoes changes in gene expression and morphology to become specialized and acquire distinct functions at a specific anatomical or pathological location <AtLoc>.
@@
Remodeling Indicated by the given trigger <Trigger>, a specific tissue <Theme> undergoes structural changes, reorganization, and modification in response to various stimuli or during growth and development.
@@
Reproduction Indicated by the given trigger <Trigger>, a specific organism <Theme> produces offspring through sexual or asexual means, leading to the continuation of the species.
@@
Mutation indicated by the given trigger <Trigger>, a specific gene, genome, or protein <Theme> undergoes a heritable change in its genetic sequence or structure at a particular site <Site>, potentially leading to alterations in its function or expression within a specific anatomical or pathological context <AtLoc>.
@@
Carcinogenesis Indicated by the given trigger <Trigger>, a specific anatomical or pathological entity <Theme> undergoes a series of genetic and cellular changes at a specific anatomical or pathological location <AtLoc>, leading to the development of cancer.
@@
Cell Transformation Indicated by the given trigger <Trigger>, a specific cell <Theme> undergoes changes in its phenotype, function, or behavior at a specific anatomical or pathological location <AtLoc>, often associated with the acquisition of abnormal or cancerous characteristics.
@@
Metastasis Indicated by the given trigger <Trigger>, a specific anatomical or pathological entity <Theme> spreads and establishes secondary growths or lesions at a different anatomical or pathological location <ToLoc> from the primary tumor site.
@@
Infection Indicated by the given trigger <Trigger>, an organism <Participant> invades and establishes itself in a specific anatomical or pathological site <Theme>, leading to a disease condition.
@@
Metabolism Indicated by the given trigger <Trigger>, collective chemical reactions occur within an organism <Theme> involving the processing, transformation, and utilization of specific molecules to maintain cellular functions and energy requirements.
@@
Synthesis Indicated by the given trigger <Trigger>, a specific simple chemical <Theme> is produced or created through chemical reactions or biological processes.
@@
Catabolism Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes chemical reactions or metabolic pathways to break down into simpler compounds, releasing energy in the process.
@@
Amino Acid Catabolism Indicated by the given trigger <Trigger>, a specific amino acid <Theme> is broken down through metabolic pathways, leading to the release of energy and the generation of byproducts like ammonia and carbon compounds.
@@
Glycolysis Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes a series of enzymatic reactions, ultimately converting glucose into pyruvate and producing ATP and NADH as energy carriers.
@@
Glycosylation Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of carbohydrate molecules (glycans) to specific sites, typically on proteins or lipids, leading to the formation of glycoproteins or glycolipids with diverse biological functions.
@@
Acetylation Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of an acetyl group at a particular site <Site>, catalyzed by another molecule <Cause>.
@@
Deacetylation Indicated by the given trigger <Trigger>, a specific molecule <Theme> has an acetyl group removed from a particular site <Site>, facilitated by another molecule <Cause>.
@@
Ubiquitination Indicated by the given trigger <Trigger>, a specific molecule <Theme> is modified by the attachment of one or more ubiquitin molecules at a particular site, facilitated by another molecule <Cause>, often involving a simple chemical group <Site> as the site of attachment.
@@
Deubiquitination Indicated by the given trigger <Trigger>, a specific molecule <Theme> has ubiquitin molecules removed from a particular site, facilitated by another molecule <Cause> involving a simple chemical group <Site> as the site of removal.
@@
Gene Expression Indicated by the given trigger <Trigger>, a specific gene, genome, or protein <Theme> is activated, leading to the production of RNA or protein and subsequent biological functions.
@@
Transcription Indicated by the given trigger <Trigger>, genetic information from a specific gene, genome, or RNA molecule <Theme> is used as a template to produce complementary RNA (usually messenger RNA) by RNA polymerase.
@@
Translation Indicated by the given trigger <Trigger>, the genetic information carried by messenger RNA <Theme> is used to synthesize a protein by ribosomes in the cell.
@@
Protein Processing Indicated by the given trigger <Trigger>, the series of post-translational modifications, folding, and transportation of a specific gene, genome, or protein <Theme> in the cell to achieve its mature and functional form.
@@
Phosphorylation Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of a phosphate group at a particular site <Site> within a protein domain or region, often regulating the molecule’s activity or function.
@@
Dephosphorylation Indicated by the given trigger <Trigger>, a specific molecule <Theme> has a phosphate group removed from a particular site <Site> within a protein domain or region, often resulting in the modulation or termination of its activity or function.
@@
DNA Methylation Indicated by the given trigger <Trigger>, a specific gene or genome <Theme> has a methyl group added to a particular site <Site> within a protein domain or region, often resulting in the regulation of gene expression and epigenetic modifications.
@@
DNA Demethylation Indicated by the given trigger <Trigger>, a specific gene or genome <Theme> has a methyl group removed from a particular site <Site> within a protein domain or region, often resulting in the modulation of gene expression and epigenetic modifications.
@@
Pathway Indicated by the given trigger <Trigger>, involving a specific molecule <Participant> that collaborates to accomplish a specific biological function or response.
@@
Binding Indicated by the given trigger <Trigger>, a specific molecule <Theme> forms a physical association or interaction with another molecule or region <Site> on a protein or DNA, potentially leading to functional changes or regulatory effects.
@@
Dissociation Indicated by the given trigger <Trigger>, a specific molecule <Theme> separates or detaches from another molecule or region <Site> on a protein or DNA, leading to the termination or disruption of their interaction or complex formation.
@@
Localization Indicated by the given trigger <Trigger>, a specific molecule <Theme> is directed to or away from a particular cellular component <AtLoc><FromLoc><ToLoc> or subcellular location within the cell.
@@
Regulation Indicated by the given trigger <Trigger>, an entity <Theme> is controlled or influenced by another entity <Cause> to achieve a specific biological effect or outcome.
@@
Positive Regulation Indicated by the given trigger <Trigger>, an entity <Theme> is promoted or enhanced by another entity <Cause> to achieve a specific biological effect or outcome.
@@
Negative Regulation Indicated by the given trigger <Trigger>, an entity <Theme> is inhibited or suppressed by another entity <Cause> to achieve a specific biological effect or outcome.
@@
Planned Process Indicated by the given trigger <Trigger>, a specific entity <Theme> is involved or manipulated using an instrument <Instrument> for a predetermined outcome or purpose.